OpenStudy (emilyisanemoji):

please help, will medal, might fan -8 + 4n = 9n + 24 what is n?

alones (alones):

So as you know start from simplify both sides of the equation \(\ 4n−8=9n+24\) and subtract \(9n\) from each sides \(\ 4n−8−9n=9n+24−9n\) Simmplified->\(−5n−8=24\)

alones (alones):

So than add \(8\) to \(−5n−8+{8}=24+8\) \(\ −5n=32\)

alones (alones):

At last divide each side by \( 5\) \[\frac{ -5n }{ 5 }=\frac{ 32 }{ -5 }\] So when you divide that your final answer should be \[n=\frac{ -32 }{ 5}\]

OpenStudy (emilyisanemoji):

thank you

alones (alones):

You'rewelcome c:'